Part ONE

Holt Enterprises recently paid a dividend, D0, of $3.00. It expects to have nonconstant growth of 20% for 2 years followed by a constant rate of 8% thereafter. The firm's required return is 19%.

How far away is the horizon date? Which Statement is Correct?

The terminal, or horizon, date is Year 0 since the value of a common stock is the present value of all future expected dividends at time zero.

The terminal, or horizon, date is the date when the growth rate becomes nonconstant. This occurs at time zero.

The terminal, or horizon, date is the date when the growth rate becomes constant. This occurs at the beginning of Year 2.

The terminal, or horizon, date is the date when the growth rate becomes constant. This occurs at the end of Year 2.

The terminal, or horizon, date is infinity since common stocks do not have a maturity date.

What is the firm's horizon, or continuing, value? Do not round intermediate calculations. Round your answer to the nearest cent. $ What is the firm's intrinsic value today, ? Do not round intermediate calculations. Round your answer to the nearest cent.

Part two

Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $0.50 coming 3 years from today. The dividend should grow rapidly - at a rate of 21% per year - during Years 4 and 5, but after Year 5, growth should be a constant 4% per year. If the required return on Computech is 12%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.

Part three

Assume today is December 31, 2019. Imagine Works Inc. just paid a dividend of $1.10 per share at the end of 2019. The dividend is expected to grow at 12% per year for 3 years, after which time it is expected to grow at a constant rate of 6% annually. The company's cost of equity (rs) is 9.5%. Using the dividend growth model (allowing for nonconstant growth), what should be the price of the company's stock today (December 31, 2019)? Do not round intermediate calculations. Round your answer to the nearest cent.
$   per share
Answer #1

value of stock = Present value of dividends + Horizontal value

Horizontal value = dividend next year/(Required return - growth rate)

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

1)

The terminal, or horizon, date is the date when the growth rate becomes constant. This occurs at the end of Year 2.

horizontal value = D3/(r-g)

= 3*1.2^2*1.08/(0.19-0.08)

= 42.4145454545

= 42.42

Intrinsic value = 3*1.2/1.19 + 3*1.2^2/1.19^2 + 42.4145454545/1.19^2

= 36.03

2)

horizontal value = 0.5*1.21^2*1.04/(0.12-0.04)

= 9.51665

value of stock = 0.5/1.12^3 + 0.5*1.21/1.12^4 + 0.5*1.21^2/1.12^5 + 9.51665/1.12^5

= 6.56

3)

horizontal value = 1.1*1.12^3*1.06/(0.095-0.06)

= 46.8041728

value of stock = 1.1*1.12/1.095 + 1.1*1.12^2/1.095^2 + 1.1*1.12^3/1.095^3 + 46.8041728/1.095^3

= 39.10